OpenStudy (anonymous):

six sandwiches are being divided evenly among 4 friends how many sandwiches will each kid recive A.2 b.3/2 c.2/3 d.3/4

OpenStudy (anonymous):

no test qns pls.

OpenStudy (anonymous):

texas chic

OpenStudy (texaschic101):

6/4 = 3/2 check.. 3/2 * 4 = 6 12/2 = 6 6 = 6 (correct) each person will receive 3/2 of a sandwich

OpenStudy (anonymous):

thx

OpenStudy (texaschic101):

anytime :)

OpenStudy (anonymous):

@texaschic101 i dont think its in the right behaviour to answer test questions man...

OpenStudy (anonymous):

shutup

OpenStudy (anonymous):

lol